Use Annual Cash Flow Analysis

  1. A warehouse can be roofed with either a painted steel material at 30 cents per square foot or galvanized steel at 50 cents per square foot. The painted roof will last 20 years; the galvanized roof will last 30 years. Assuming an APR = 6.5%, determine the most economical choice.
Answer #1

The equivalent annual cost of painted steel material per sqft = 0.3 *(A/P, 6.5%,20)

= 0.3 * 0.090756

= 0.027227

The equivalent annual cost of galvanized steel material per sqft = 0.5 *(A/P, 6.5%,30)

= 0.5 * 0.076577

= 0.038289

As the equivalent annual cost of painted steel material per sqft is less, it should be selected